A car moving along straight road at 30m/s slows uniformly to a speed of 10m/s in a time of 5.0s. Determine


(a) acceleration of the car


(b) distance it moves during the third second

1
Expert's answer
2022-02-18T15:53:31-0500

Given:

"v_1=30\:\rm m/s"

"v_2=10\:\rm m/s"

"\Delta t=5.0\:\rm s"


(a) the acceleration of the car

"a=\frac{\Delta v}{\Delta t}=\frac{10-30}{5.0}=-4\:\rm m/s^2"

(b) distance it moves during the third second

"d(3)-d(2)=(30*3-4*3^2/2)\\
-(30*2-4*2^2/2)=20\:\rm m"